PAT HAD TO RUN SOME ERRANDS THAT TOOK HER TO SEVERAL TOWNS IN HER AREA. SHE LEFT ANDERSON AND DROVE TO BANDOR AND THEN CONTINUED ON TO CARYVILLE. AFTER SHE FINISHED HER ERRANDS, SHE RETURNED TO HER HOME IN ANDERSON. THE DISTANCE TO BAND OR TO CARRYVILLE WAS 5 miles less than twice the distance from ANDERSON to band or, the distance from Caryville to Anderson was 4 miles more than the distance from ANDERSON to Band or. If the total miles she drove was 63 miles, than what was the distance from ANDERSON to Bandor? From Bandor to Caryville? From Caryville to Anderson? Where do I start?

Answer by stanbon(75887)

PAT HAD TO RUN SOME ERRANDS THAT TOOK HER TO SEVERAL TOWNS IN HER AREA. SHE LEFT ANDERSON AND DROVE TO BANDOR AND THEN CONTINUED ON TO CARYVILLE. AFTER SHE FINISHED HER ERRANDS, SHE RETURNED TO HER HOME IN ANDERSON.
----------------------------------
THE DISTANCE From BANDOR TO CARRYVILLE WAS 5 miles less than twice the distance from ANDERSON to BANDOR.
C-B = 2(B-A)-5
So, B-A = [(C-B)+5]/2
=====================
The distance from Caryville to Anderson was 4 miles more than the distance from ANDERSON to BANDOR.
A-C = B-A +4
So, B-A = (A-C)-4
=====================
If the total miles she drove was 63 miles
(B-A) + (C-B) + (A-C) = 63
2(B-A) - (C-B)+ 0 = 5
(B-A) + 0 - (A-C) = -4
--------------------------------
Solve the system of equations:
I get:
B-A = 16 miles
C-A = 27 miles
A-C = 20 miles
--------
What was the distance from ANDERSON to Bandor? 16
From Bandor to Caryville? 27
From Caryville to Anderson? 20
====================
Cheers,
Stan H.
====================